OpenStudy (anonymous):

BEST RESPONSE Find a ⋅ b. a = <5, 2>, b = <4, 5>

OpenStudy (anonymous):

multiply the two x terms together and multiply the two y terms together. put the results into another ordered pair.

OpenStudy (anonymous):

okay can you explain how to do this one? Find a ⋅ b. a = 10i + 9j, b = 4i + 3j is it the same process? @tomo

OpenStudy (anonymous):

is this two vectors?

OpenStudy (anonymous):

dot products of two vectors yes

OpenStudy (anonymous):

@tomo I don't believe that is correct. If I am right, this is the dot product: if x = <a,b> and y = <c,d>, the dot product is a number!! whose value is ac + bd In your case, the dot product is 30.

OpenStudy (anonymous):

yes you are correct @khoala4pham

OpenStudy (anonymous):

so how would i do a = 10i + 9j, b = 4i + 3j @khoala4pham

OpenStudy (anonymous):

then you have <10,9> <4,3> just follow the same process

OpenStudy (anonymous):

okay i have one more quick question!

OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (anonymous):

The same is true for n-dimensional vectors: you take the product of each corresponding coordinate and then take the sum of all of it. Generally, <a1,a2,a3,a4,...an> (dot) <b1,b2,b3,b4,...bn> = a1*b1 + a2*b2 + a3*b3 + a4*b4 +...(an)*(bn) In your case, v (dot) w is 2(-2) + (-8)6 + (-8)(-5) = -4 - 48 + 40= -12

OpenStudy (anonymous):

Thank you!